1 If f(x) = 4x + 5, what is the value of f(2)? A-2 B 13 Drag the red circle to choose your answer. C 17 D4 O
Answer #1

f(1) = =unts Put a 2, we get fra) f(2)=4 (2) + 5 = 8+5 f(a)=13 Option (B) is correct answer